OpenStudy (bernalcheryl):

Help me solve this problem please 2.1x-1.34=0.42-6.7x

OpenStudy (anonymous):

x=.2

OpenStudy (bernalcheryl):

and how did you get the answer?

OpenStudy (anonymous):

2.1x - 1.34 = 0.42 - 6.7x subtract -6.7x from both sides: 2.1x - 1.34 - (-6.7x) = 0.42 - 6.7x - (-6.7x) => 8.8x - 1.34 = 0.42 subtract -1.34 from both sides: 8.8x - 1.34 - (-1.34) = 0.42 - (-1.34) => 8.8x = 1.76 divide both sides by 8.8: 8.8x/8.8 = 1.76/8.8 => **x = 0.2**